# Autodeleveraging: Impossibilities and Optimization

### v1

@@ 2 Background
Note this toy example is net short by 2 contracts: $\sum_{i}b_{i}q_{i}=-2$. For AMM-style venues, the pool holds the offsetting inventory $q_{V}=2$ with $b_{V}=+1$, so across traders plus venue $\sum_{i}b_{i}q_{i}+b_{V}q_{V}=0$. For order-book venues, interpret $\mathcal{P}_{5}$ as a subset of accounts. The complementary matched positions are omitted but implied. [^5] We note that in centralized exchanges, the exchange itself creates the complementary positions whereas in decentralized exchanges with order books (like Hyperliquid and Lighter), a liquidity provider like HLP or LLP owns the position; see Hyperliquid’s HLP documentation [HyperliquidHLPVaults] and the Lighter whitepaper [LighterWhitepaper]. All zero-sum statements (e.g., funding rates, which will be defined shortly) are taken over traders plus the venue inventory.
@@
The latter condition is needed because the perpetuals exchange is meant to be a neutral matching layer for traders. This means that the exchange should not have a net exposure to the spot asset, as it increases the risk of default for traders. Specifically, the position $\overline{\mathfrak{p}}$ is created to ensure that the total notional exposure of the exchange to the spot asset is zero. [^6] In centralized venues such as Binance Futures, the matching engine and clearing layer are designed to operate a matched book over user positions, as reflected in their open interest endpoints for USD\$- and COIN-margined futures [BinanceFuturesOpenInterest]. In decentralized order-book venues like Hyperliquid, the protocol core similarly avoids taking directional exposure. Offsetting inventory is held in separate HLP vaults on behalf of liquidity providers [HyperliquidHLPVaults]. We define
@@
Order book venues maintain a limit order book of bids and asks for perpetual futures contracts at discrete price levels, and have been extensively studied in the market microstructure literature [Kyle1985, AlmgrenChriss2001, Gatheral2010, BouchaudImpact2010]. Traders may submit limit orders, which rest in the book until they are matched, or market orders, which execute immediately against resting liquidity. When a trader opens a new position as a taker, the execution price $p_{t}$ is the volume-weighted average of the resting orders they consume, i.e., the price they would obtain by submitting a market order of size $q$ into the book at time $t$. The largest order book venues in the perpetual futures market include centralized exchanges such as Binance and Bybit and decentralized exchanges such as Hyperliquid and Lighter. [^7] From a risk-warehousing perspective, Hyperliquid’s HLP vaults cause the protocol to function as a hybrid between a pure order-book venue and an inventory-taking pool: user orders are matched via an order book, but directional exposure is warehoused collectively in HLP vaults on behalf of liquidity providers [HyperliquidHLPVaults, OakResearchJELLY, CoinDesk2025LargestLiquidations].
Automated market maker (AMM) venues maintain a pool of collateral and inventory for the perpetual, and quote prices according to a deterministic pricing rule or bonding curve that depends on the pool state, typically implemented via constant-function market makers [AngerisChitra2020ImprovedOracles, Angeris2023GeometryCFMM]. Traders interact with the pool via swap-like trades: submitting an order of size $q$ moves the pool along its pricing curve, and the execution price $p_{t}$ is the average price paid over this path. When a trader opens a new position against the AMM, the pool itself holds the offsetting inventory $\overline{\mathfrak{p}}$, so that the trader’s exposure is matched by an equal and opposite position in the pool, as in inventory-taking perpetual venues such as GMX, Drift, and Perpetual Protocol v2 [GMXDocs, DriftADLCode, PerpV2Docs].
@@
Most funding rates are simple linear functions of the relative imbalance between the long and short positions. A common funding rate [Chitra2025PDLP] formula takes the form:
@@ 2.2 Liquidations
When a user’s position is insolvent, i.e., (7) holds, the exchange must liquidate the position — that is, remove some or all of the position from $\mathcal{P}_{n}$. Liquidations are used to ensure that the exchange itself remains solvent, which ensures that traders can withdraw their funds and profits as expected. The process of liquidating a position is probabilistic, however, and can fail in a number of ways. We will illustrate the high-level process of liquidation using bankruptcy and liquidation prices. Then we will describe how exchanges can lose solvency when liquidations fail, leading to the usage of autodeleveraging mechanisms. Our description of liquidations is minimal, avoiding formal mathematical description unless needed, as we are concerned with liquidations failure versus liquidation mechanics. We refer the interested reader to [Soska2021BitMEX, perez2021liquidations, qin2021empirical, AngerisChitra2023PerpsSIAM] for detailed accounts of liquidation mechanisms and modeling on centralized and decentralized venues.
@@ 2.2.1 Liquidation Prices
Before defining how liquidations work, we need to define the criteria for when a position is eligible for liquidation. These criteria depend on when certain prices are reached where a trader position has low or negative equity. We focus on defining three prices: bankruptcy price (zero equity), liquidation price (low equity), and execution price (actual realized price of a liquidation). In practice, liquidation mechanisms often use more price-based variables to decide on execution. However, all mechanisms define these three prices, allowing us to abstract liquidations to mechanisms involving these prices.
@@
This formula represents the common maxim that a perpetuals position with $\ell_{i}\geq 1$ times leverage will be liquidated when the price moves by $\frac{1}{\ell_{i}}$% in the direction opposite to $b_{i}$; see, for example, [Binance2025CollateralLeverageUpdate, he2022fundamentals]. For instance, a maximal 10x leveraged position will be liquidated when the price moves by 10% from the initial price. Detailed numerical examples of bankruptcy prices are provided in Appendix A.
@@
Detailed numerical examples of liquidation prices are provided in Appendix A.
@@
Shortfalls are first absorbed by the insurance fund (§2.3) up to its balance. Any residual shortfall not covered by the insurance fund is socialized via ADL (§2.4). The choice of liquidation size $\Delta q$ and impact parameter $\alpha$ together determines whether thin-equity positions can typically be closed without realizing bad debt. Detailed numerical examples of execution prices and shortfalls are provided in Appendix A.
@@ 2.2.2 Liquidation Mechanics
Given liquidation prices, a natural question is how liquidations are mechanically executed. At a high level, the exchange seizes a low equity position’s cash and collateral and sells it to the market. Such a sale can realize a profit or a loss for the exchange and is best thought of as a trading strategy itself. We will provide a high-level description of liquidation mechanisms as trading strategies, with the caveat that there are many idiosyncrasies in practical implementations (see, e.g., [BinanceADL, HyperliquidDocsLiquidations, BitMEXADL] for venue documentation and differences between auctions, order-book liquidations, and RFQ-style closures).
@@
In decentralized exchanges, liquidators are usually any market participant with enough capital who can connect to the exchange. On the other hand, in centralized exchanges, liquidators are usually whitelisted parties approved by the exchange to perform liquidations We denote the set of liquidation costs for a liquidation of size $\Delta q$ at time $t$ as $\tau_{t}(\Delta q)\in{\mbox{\bf R}}_{+}$, which represents the cash cost paid by the user upon liquidation. Detailed numerical examples of liquidation costs are provided in Appendix A.
@@
The greedy liquidation strategy simply takes the smallest feasible root of this quadratic and uses it as the liquidation quantity $\Delta q$. We refer the reader to Appendix A for detailed numerical examples of this strategy and related liquidation procedures.
@@
where $\Delta q_{i}=L(\mathfrak{p}_{i,t},p_{1:T},\hat{p}_{1:T})$. This represents the shortfall that the exchange must cover to be solvent (§2.3). Detailed numerical examples of bad debt are provided in Appendix A.
@@
Given the bankruptcy, liquidation, and execution prices, we can now describe the high-level algorithm that liquidations follow (see Appendix A). We note that many live liquidation systems will have much more complex liquidation algorithms. These complexities deal with the coordination costs of coordinating many parties (e.g. oracle provider, liquidators, spot order book liquidity) and precise models that exchanges use for their liquidation strategy. However, we effectively lump all of these complexities into the definition of the liquidation strategy. A detailed algorithmic description of the liquidation loop and a step-by-step example are provided in Appendix A.
@@ 2.3 Exchange Solvency
An exchange is said to be insolvent if the following condition holds $\mathsf{Solv}_{T}(\mathcal{P}_{n})\leq 0$. The insolvency condition corresponds to when the total cash held by the exchange on behalf of users is less than the total accrued profits. This mean that if this insolvency condition is hit, the exchange will not be about to payout all profitable traders when they withdraw from the exchange. Note that this is a global condition on $\mathcal{P}_{n}$ as opposed to a per user constraint.
@@
Insurance reserves in crypto funds were first popularized alongside perpetual futures in the mid-2010s (notably on BitMEX) [Soska2021BitMEX]. A majority of perpetuals venues (centralized and on-chain) maintain such reserves (e.g. BitMEX [BitMEXADL], Binance [BinanceFuturesInsuranceFund], Drift [DriftADLCode], Bybit [BybitInsuranceFund]). To illustrate the size of these funds, we note that mid-2025 estimates place OKX and Bybit insurance funds at roughly \$300M and \$150M, respectively [OKXBybitInsuranceComparison].
@@
We note that when $R_{t}>0$, an ADL mechanism is triggered.
@@
A natural question to ask is, what is the optimal size for an insurance fund? While this question is nuanced in practice, we provide a simple mean-field model that provide some intuition. In particular, we determine the optimal fund size $K^{*}$ by minimizing a total cost function that balances the opportunity cost of capital against the expected franchise damage from ADL events. Let $r>0$ denote the opportunity cost of holding capital (e.g., the risk-free rate) and $\kappa>0$ denote the reputation cost per unit of socialized loss (reflecting lost future volume). The exchange minimizes
@@
This is a classic newsvendor-type problem [ArrowHarrisMarschak1951]. Classical result show that the optimal size is the Value-at-Risk of the deficit distribution at quantile $1-r/\kappa$:
@@
For completeness, we provide a derivation in Appendix A.8. Intuitively, as the reputation cost $\kappa$ increases relative to the cost of capital $r$, the exchange holds a larger buffer to push ADL events further into the tail.
@@ 2.4 Autodeleveraging
We first note that ADL mechanisms are not deterministic and foolproof. For example, if an exchange is already insolvent, $\mathsf{Solv}_{T}(\mathcal{P}_{n})\leq 0$, then no deterministic ADL mechanism can return solvency. To see this, we define four key quantifies, the total shortfall $D_{T}(\mathcal{P}_{n})$, the maximum shortfall $\Delta_{T}(\mathcal{P}_{n})$, the total winner’s equity $W_{T}(\mathcal{P}_{n})$, and the max-winner $\omega_{T}(\mathcal{P}_{n})$:
$D_{T}(\mathcal{P}_{n})$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}\big(-e_{T}(\mathfrak{p})\big)_{+}\geq 0$ $\Delta_{T}(\mathcal{P}_{n})$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}\big(-e_{T}(\mathfrak{p})_{-}\big)\geq 0$
$W_{T}(\mathcal{P}_{n})$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}e_{T}(\mathfrak{p})_{+}\geq 0$ $\omega_{T}(\mathcal{P}_{n})$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}e_{T}(\mathfrak{p})_{+}\geq 0$
For notational simplicity, we will use Roman letters for total quantities and Greek letters for maximum or extreme value quantities. The shortfall represents the total amount of negative equity whereas the winners’ equity represents the total amount of positive equity. By definition, $\mathsf{Solv}_{T}(\mathcal{P}_{n})=W_{T}(\mathcal{P}_{n})-D_{T}(\mathcal{P}_{n})$, so if $\mathsf{Solv}_{T}(\mathcal{P}_{n})\leq 0$, then $D_{T}(\mathcal{P}_{n})\geq W_{T}(\mathcal{P}_{n})$, i.e., no amount of forced liquidations of positive equity can cover the shortfall.
@@
An ADL policy $\pi(\mathcal{P}_{n})$ maps a perpetuals exchange to a fraction of shortfall to socialize, $\theta\in[0,1]$, and a set of haircuts to apply to positions, $h\in[0,1]^{n}$. We term the fraction $\theta$ the severity of the ADL policy. For notational convenience we define $\theta_{\pi},h_{\pi}$ as the severity and haircuts induced by an ADL policy $\pi$. We say an ADL policy $\pi$ if valid if the following three constraints always hold:
$\sum_{\mathfrak{p}\in\mathcal{P}_{n}}h_{\pi,i}e_{T}(\mathfrak{p})_{+}=\theta_{\pi}D_{T}(\mathcal{P}_{n})=-\sum_{\mathfrak{p}\in\mathcal{P}_{n}}h_{\pi,i}e_{T}(\mathfrak{p})_{-}.$ (17)
$\theta_{\pi}D_{T}(\mathcal{P}_{n})\leq W_{T}(\mathcal{P}_{n}).$ (18)
where $p_{i}$ is the $i$th trader’s position. The first two conditions (17) ensure that when an ADL policy is executed, only a $\theta$% of the deficit is socialized from the winners to the losers (i.e., a budget balance constraint). The third condition (18) ensures that $\theta_{\pi}$ is chosen such that one does not need to haircut more equity than what the winners have (i.e., a feasibility constraint). We define the post-policy deficit $D^{\pi}_{T}$, winners $W^{\pi}_{T}$, and max-winner $M^{\pi}_{T}$ as
$D^{\pi}_{T}$ $=\theta_{\pi}D_{T}(\mathcal{P}_{n})$ $\Delta^{\pi}_{T}$ $=\theta_{\pi}\Delta_{T}(\mathcal{P}_{n})$
$W^{\pi}_{T}$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}(1-h_{\pi,i})(e_{T}(\mathfrak{p}))_{+}$ $\omega^{\pi}_{T}$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}(1-h_{\pi,i})(e_{T}(\mathfrak{p}))_{+}$
@@
Given an ADL policy outputting $(\theta_{\pi},h_{\pi})$, the positive equity positions $\mathcal{W}_{T}=\{i:e_{T}(\mathfrak{p}_{i})>0\}$ are haircut by $h_{\pi,i}$, leading to equities $e_{T}(\mathfrak{p}_{i})\leftarrow(1-h_{\pi,i})e_{T}(\mathfrak{p}_{i})$ for $i\in\mathcal{W}_{T}$. For instance, if we have equities $e_{T}=(10,5,-5)$ and the ADL policy outputs haircuts (0.5,0,1)(0.5,0,-1), then the post-ADL equities are $e_{T}=(5,5,0)$. On the other hand, if the ADL policy outputs haircuts h=(0.25,0.5,1)h=(0.25,0.5,-1), then the final equities after execution of the policy are $e_{T}=(7.5,2.5,0)$.
@@
PNL and Leverage Ranking.
The policy that the largest centralized exchange, Binance, and the largest decentralized exchange, Hyperliquid, use is the PNL-leverage ranking (which we also refer to as a queueing method). We will denote this strategy as $\pi_{PL}$ and define it via the algorithm that implements it. Given a haircut budget $B=\theta_{\pi_{PL}}D_{T}(\mathcal{P}_{n})$, the algorithm to choose $h_{i}$ works as follows:
@@
    For each position $\mathfrak{p}=(q,c,\ell,t,b)\in\mathcal{P}_{n}$, define the score [^9] We note that technically, many exchanges define the leverage used in the score differently (i.e., instead of the score being linear in $\ell$, some exchanges multiply by $\frac{\ell pq}{pq+c})$. This does not change our main results and mainly adds technical complications to the example. $s_{T}(\mathfrak{p},p^{ref})=\ell\,\frac{\hat{p}_{T}}{p^{ref}}$.

### v3

@@ 2 Background
Note this toy example is net short by 2 contracts: $\sum_{i}b_{i}q_{i}=-2$. For AMM-style venues, the pool holds the offsetting inventory $q_{V}=2$ with $b_{V}=+1$, so across traders plus venue $\sum_{i}b_{i}q_{i}+b_{V}q_{V}=0$. For order-book venues, interpret $\mathcal{P}_{5}$ as a subset of accounts. The complementary matched positions are omitted but implied. [^5] We note that in centralized exchanges, the exchange itself creates the complementary positions whereas in decentralized exchanges with order books (like Hyperliquid and Lighter), a liquidity provider like HLP or LLP owns the position; see Hyperliquid’s HLP documentation [71] and the Lighter whitepaper [82]. All zero-sum statements (e.g., funding rates, which will be defined shortly) are taken over traders plus the venue inventory.
@@
The latter condition is needed because the perpetuals exchange is meant to be a neutral matching layer for traders. This means that the exchange should not have a net exposure to the spot asset, as it increases the risk of default for traders. Specifically, the position $\overline{\mathfrak{p}}$ is created to ensure that the total notional exposure of the exchange to the spot asset is zero. [^6] In centralized venues such as Binance Futures, the matching engine and clearing layer are designed to operate a matched book over user positions, as reflected in their open interest endpoints for USD\$- and COIN-margined futures [24]. In decentralized order-book venues like Hyperliquid, the protocol core similarly avoids taking directional exposure. Offsetting inventory is held in separate HLP vaults on behalf of liquidity providers [71]. We define
@@
Order book venues maintain a limit order book of bids and asks for perpetual futures contracts at discrete price levels, and have been extensively studied in the market microstructure literature [79, 8, 58, 27]. Traders may submit limit orders, which rest in the book until they are matched, or market orders, which execute immediately against resting liquidity. When a trader opens a new position as a taker, the execution price $p_{t}$ is the volume-weighted average of the resting orders they consume, i.e., the price they would obtain by submitting a market order of size $q$ into the book at time $t$. The largest order book venues in the perpetual futures market include centralized exchanges such as Binance and Bybit and decentralized exchanges such as Hyperliquid and Lighter. [^7] From a risk-warehousing perspective, Hyperliquid’s HLP vaults cause the protocol to function as a hybrid between a pure order-book venue and an inventory-taking pool: user orders are matched via an order book, but directional exposure is warehoused collectively in HLP vaults on behalf of liquidity providers [71, 89, 99].
Automated market maker (AMM) venues maintain a pool of collateral and inventory for the perpetual, and quote prices according to a deterministic pricing rule or bonding curve that depends on the pool state, typically implemented via constant-function market makers [10, 11]. Traders interact with the pool via swap-like trades: submitting an order of size $q$ moves the pool along its pricing curve, and the execution price $p_{t}$ is the average price paid over this path. When a trader opens a new position against the AMM, the pool itself holds the offsetting inventory $\overline{\mathfrak{p}}$, so that the trader’s exposure is matched by an equal and opposite position in the pool, as in inventory-taking perpetual venues such as GMX, Drift, and Perpetual Protocol v2 [60, 47, 94].
@@
Most funding rates are simple linear functions of the relative imbalance between the long and short positions. A common funding rate [33] formula takes the form:
@@ 2.2 Liquidations
When a user’s position is insolvent, i.e., (7) holds, the exchange must liquidate the position — that is, remove some or all of the position from $\mathcal{P}_{n}$. Liquidations are used to ensure that the exchange itself remains solvent, which ensures that traders can withdraw their funds and profits as expected. The process of liquidating a position is probabilistic, however, and can fail in a number of ways.
We will illustrate the high-level process of liquidation using bankruptcy and liquidation prices. Then we will describe how exchanges can lose solvency when liquidations fail, leading to the usage of autodeleveraging mechanisms. Our description of liquidations is minimal, avoiding formal mathematical description unless needed, as we are concerned with liquidation failure versus liquidation mechanics. We refer the interested reader to [107, 93, 98, 9] for detailed accounts of liquidation mechanisms and modeling on centralized and decentralized venues.
@@ 2.2.1 Liquidation Prices
Before defining how liquidations work, we need to define criteria for when a position is eligible for liquidation. These criteria depend on when certain prices are reached where a trader position has low or negative equity. We focus on defining three prices: bankruptcy price (zero equity), liquidation price (low equity), and execution price (actual realized price of a liquidation). In practice, liquidation mechanisms often use more price-based variables to decide on execution. However, all mechanisms define these three prices, allowing us to abstract liquidations to mechanisms involving these prices.
@@
This formula represents the common maxim that a perpetuals position with $\ell_{i}\geq 1$ times leverage will be liquidated when the price moves by $\frac{1}{\ell_{i}}$% in the direction opposite to $b_{i}$; see, for example, [23, 67]. For instance, a maximal 10x leveraged position will be liquidated when the price moves by 10% from the initial price. Detailed numerical examples of bankruptcy prices are provided in Appendix B.
@@
Detailed numerical examples of liquidation prices are provided in Appendix B.
@@
Shortfalls are first absorbed by the insurance fund (§2.3) up to its balance. Any residual shortfall not covered by the insurance fund is socialized via ADL (§2.4). The choice of liquidation size $\Delta q$ and impact parameter $\alpha$ together determines whether thin-equity positions can typically be closed without realizing bad debt. Detailed numerical examples of execution prices and shortfalls are provided in Appendix B.
@@ 2.2.2 Liquidation Mechanics
Given liquidation prices, a natural question is how liquidations are mechanically executed. At a high level, the exchange seizes a low equity position’s cash and collateral and sells it to the market. Such a sale can realize a profit or a loss for the exchange and is best thought of as a trading strategy itself. We will provide a high-level description of liquidation mechanisms as trading strategies, with the caveat that there are many idiosyncrasies in practical implementations (see, e.g. [21, 69, 25] for venue documentation and differences between auctions, order-book liquidations, and RFQ-style closures).
@@
In decentralized exchanges, liquidators are usually any market participant with enough capital who can connect to the exchange. On the other hand, in centralized exchanges, liquidators are usually whitelisted parties approved by the exchange to perform liquidations We denote the set of liquidation costs for a liquidation of size $\Delta q$ at time $t$ as $\tau_{t}(\Delta q)\in{\mbox{\bf R}}_{+}$, which represents the cash cost paid by the user upon liquidation. Detailed numerical examples of liquidation costs are provided in Appendix B.
@@
The greedy liquidation strategy simply takes the smallest feasible root of this quadratic and uses it as the liquidation quantity $\Delta q$. We refer the reader to Appendix B for detailed numerical examples of this strategy and related liquidation procedures.
@@
where $\Delta q_{i}=L(\mathfrak{p}_{i,t},p_{1:T},\hat{p}_{1:T})$. This represents the shortfall that the exchange must cover to be solvent (§2.3). Detailed numerical examples of bad debt are provided in Appendix B.
@@
Given the bankruptcy, liquidation, and execution prices, we can now describe the high-level algorithm that liquidations follow (see Appendix B). We note that many live liquidation systems will have much more complex liquidation algorithms. These complexities deal with the coordination costs of coordinating many parties (e.g. oracle provider, liquidators, spot order book liquidity) and precise models that exchanges use for their liquidation strategy. However, we effectively lump all of these complexities into the definition of the liquidation strategy. A detailed algorithmic description of the liquidation loop and a step-by-step example are provided in Appendix B.
@@ 2.3 Exchange Solvency
An exchange is said to be insolvent if the following condition holds $\mathsf{Solv}_{T}(\mathcal{P}_{n})\leq 0$. The insolvency condition corresponds to when the total cash held by the exchange on behalf of users is less than the total accrued profits. This mean that if this insolvency condition is hit, the exchange will not be able to payout all profitable traders when they withdraw from the exchange. Note that this is a global condition on $\mathcal{P}_{n}$ as opposed to a per user constraint.
@@
Insurance reserves in crypto funds were first popularized alongside perpetual futures in the mid-2010s (notably on BitMEX) [107]. A majority of perpetuals venues (centralized and on-chain) maintain such reserves (e.g. BitMEX [25], Binance [22], Drift [47], Bybit [31]). To illustrate the size of these funds, we note that mid-2025 estimates place OKX and Bybit insurance funds at roughly \$300M and \$150M, respectively [73].
@@
We note that when $R_{t}>0$, an ADL mechanism is triggered. We term times $t$ such that $R_{t}>0$ as breach times. The frequency of breach events is referred to as breach frequency or breach rate.
@@
A natural question to ask is, what is the optimal size for an insurance fund? While this question is nuanced in practice, we provide a simple mean-field model that provides some intuition. In particular, we determine the optimal fund size $K^{*}$ by minimizing a total cost function that balances the opportunity cost of capital against the expected franchise damage from ADL events. Let $r>0$ denote the opportunity cost of holding capital (e.g. the risk-free rate) and $\kappa>0$ denote the reputation cost per unit of socialized loss (reflecting lost future volume). The exchange minimizes
@@
This is a classic newsvendor-type problem [12]. Classical result show that the optimal size is the Value-at-Risk of the deficit distribution at quantile $1-r/\kappa$:
@@
For completeness, we provide a derivation in Appendix B.8. Intuitively, as the reputation cost $\kappa$ increases relative to the cost of capital $r$, the exchange holds a larger buffer to push ADL events further into the tail.
@@ 2.4 Autodeleveraging
Contract space vs wealth space.
Production ADL is executed in contract space: the engine selects positions by a ranking score and forces contract-level reductions. The theoretical analysis in this paper is written in wealth space (equity haircuts and haircutable endowment). To compare production ADL to model ADL, we need an explicit mapping from contract closures to wealth-space impacts. This mapping is non-trivial: a close in contracts is not the same statement as haircutting 100% of equity, because the economic effect depends on prices and on how much of the account’s equity is attributable to the seized contracts. Production ADL induces wealth-space haircuts only through the observation-model mapping (see §9). It is critical to distinguish between:
  • Wealth-space queue (theory object): a greedy allocation rule that ranks positions by a score and allocates haircuts in endowment space until the budget is met.
  • Contract-space queue (production): Hyperliquid’s implementation ranks positions and closes contracts; this induces wealth-space haircuts only through the observation-model mapping (two-pass replay comparing ADL-on vs ADL-off equity outcomes).
  • Wealth-pro-rata (theory): each winner loses the same fraction of their haircutable endowment, defined in (24).
  • Contracts-pro-rata (production): some exchanges allocate losses proportional to contract size; this is a different mechanism with different fairness properties.
Without this separation, readers conflate wealth-space theory results with contract-space implementation and misread empirical overshoot as a production ledger claim. These are different objects: the theory queue is an abstraction that shares the same ranking idea as production, but does not reproduce contract-level execution details.
Equity decomposition and haircut numéraire.
We decompose each position’s equity into protected cash/principal and profit-and-loss:
$e_{i,t}=c_{i,t}+\mathrm{PNL}_{i,t},$ (17)
where $c_{i,t}\geq 0$ is the protected cash/principal (returned on forced close; not haircutable in production) and $\mathrm{PNL}_{i,t}$ is the profit-and-loss. We define the haircutable endowment $w_{i,t}\geq 0$ as the only quantity ADL may seize from position $i$ at time $t$. In the corrected measurement model used throughout this paper, we specialize to PNL-only haircuts:
$w_{i,t}=(\mathrm{PNL}_{i,t})_{+}.$ (18)
This reflects the production reality that principal/collateral is returned when a position is force-closed; counting it as “haircut capacity” would be a numéraire error. (For theoretical comparison, one may also consider an equity-haircut abstraction $w_{i,t}=(e_{i,t})_{+}$, but this is explicitly labeled as a theoretical simplification, not production.)
Deficit and capacity definitions.
We define four key quantities. The total shortfall (deficit) and maximum shortfall are computed from loser-side negative equity:
$D_{T}(\mathcal{P}_{n})$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}\big(-e_{T}(\mathfrak{p})\big)_{+}\geq 0,$ $\Delta_{T}(\mathcal{P}_{n})$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}\big(-e_{T}(\mathfrak{p})_{-}\big)\geq 0.$
The total haircut capacity and maximum haircut capacity are computed from the haircutable endowment:
$U_{T}(\mathcal{P}_{n})$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}w_{T}(\mathfrak{p})\geq 0,$ $\upsilon_{T}(\mathcal{P}_{n})$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}w_{T}(\mathfrak{p})\geq 0.$
For notational simplicity, we will use Roman letters for total quantities and Greek letters for maximum or extreme value quantities. The shortfall represents the total amount of negative equity (loser-side), whereas the capacity represents the total amount of haircutable endowment available from winners. By definition, $\mathsf{Solv}_{T}(\mathcal{P}_{n})=\sum_{\mathfrak{p}}e_{T}(\mathfrak{p})=W_{T}(\mathcal{P}_{n})-D_{T}(\mathcal{P}_{n})$, where $W_{T}(\mathcal{P}_{n})=\sum_{\mathfrak{p}}(e_{T}(\mathfrak{p}))_{+}$ is the total winner equity. However, under PNL-only haircuts, solvency ($\mathsf{Solv}_{T}>0$) does not imply the deficit is coverable by ADL. The correct one-round feasibility condition is $\theta_{T}D_{T}\leq U_{T}$ (not $\theta_{T}D_{T}\leq W_{T}$). There is an intermediate regime where $D_{T}<W_{T}$ but $D_{T}>U_{T}$: the exchange is solvent in gross equity but cannot clear the deficit without touching principal.
@@
An ADL policy $\pi(\mathcal{P}_{n})$ maps a perpetuals exchange to a fraction of shortfall to socialize, $\theta\in[0,1]$, and a set of seized amounts $x_{\pi,i}\in[0,w_{i,T}]$ (or equivalently, haircut fractions $h_{\pi,i}\in[0,1]$ such that $x_{\pi,i}=h_{\pi,i}w_{i,T}$). We term the fraction $\theta$ the severity of the ADL policy. For notational convenience we define $\theta_{\pi},h_{\pi}$ as the severity and haircuts induced by an ADL policy $\pi$. We say an ADL policy $\pi$ is valid if the following constraints always hold:
$\sum_{\mathfrak{p}\in\mathcal{P}_{n}}h_{\pi,i}w_{T}(\mathfrak{p})=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}x_{\pi,i}=\theta_{\pi}D_{T}(\mathcal{P}_{n}).$ (19)
$\theta_{\pi}D_{T}(\mathcal{P}_{n})\leq U_{T}(\mathcal{P}_{n})=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}w_{T}(\mathfrak{p}).$ (20)
The budget balance constraint (19) ensures that when an ADL policy is executed, exactly $\theta$% of the deficit is socialized from the winners’ haircutable endowment. The feasibility constraint (20) ensures that $\theta_{\pi}$ is chosen such that the required seizure does not exceed the total haircut capacity $U_{T}$. Note that under PNL-only haircuts ($w=(\mathrm{PNL})_{+}$), feasibility is stricter than solvency: even if $\mathsf{Solv}_{T}=W_{T}-D_{T}>0$, we may have $D_{T}>U_{T}$, meaning the deficit cannot be covered without touching principal.
We define the post-policy deficit $D^{\pi}_{T}$, capacity $U^{\pi}_{T}$, and max-winner endowment $\upsilon^{\pi}_{T}$ as
$D^{\pi}_{T}$ $=\theta_{\pi}D_{T}(\mathcal{P}_{n}),$ $\Delta^{\pi}_{T}$ $=\theta_{\pi}\Delta_{T}(\mathcal{P}_{n}),$
$U^{\pi}_{T}$ $=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}(w_{T}(\mathfrak{p})-x_{\pi,i})_{+}=\sum_{\mathfrak{p}\in\mathcal{P}_{n}}(1-h_{\pi,i})w_{T}(\mathfrak{p}),$ $\upsilon^{\pi}_{T}$ $=\max_{\mathfrak{p}\in\mathcal{P}_{n}}(w_{T}(\mathfrak{p})-x_{\pi,i})_{+}.$
@@
Given an ADL policy outputting $(\theta_{\pi},h_{\pi})$ (or equivalently, seized amounts $x_{\pi,i}$), the post-ADL equity for winners is:
$e^{\prime}_{i,T}=c_{i,T}+\mathrm{PNL}_{i,T}-x_{\pi,i}=c_{i,T}+\mathrm{PNL}_{i,T}-h_{\pi,i}w_{i,T}.$ (21)
For PNL-only haircuts ($w_{i,T}=(\mathrm{PNL}_{i,T})_{+}$), this becomes:
$e^{\prime}_{i,T}=c_{i,T}+\mathrm{PNL}_{i,T}-h_{\pi,i}(\mathrm{PNL}_{i,T})_{+}.$ (22)
For instance, if we have equities $e_{T}=(10,5,1,-3,-12)$ with cash components $c_{T}=(2,1,0.5,0,0)$ and PNL components $\mathrm{PNL}_{T}=(8,4,0.5,-3,-12)$, then under PNL-only haircuts with $w_{T}=(8,4,0.5,0,0)$ and policy outputs $h=(0.5,1,0,0,0)$ (seizing $x=(4,4,0,0,0)$), the post-ADL equities are $e^{\prime}_{T}=(2+8-4,1+4-4,0.5+0.5-0,-3,-12)=(6,1,1,-3,-12)$.
@@
Queue.
The policy that the largest centralized exchange, Binance, and the largest decentralized exchange, Hyperliquid, use is the Queue policy. The implementation at these exchanges uses the PNL-leverage ranking. However, the Queue model is somewhat more generic, as we define it below.
At a high level, the Queue strategy is a greedy algorithm that ranks positions using a score function, $s_{T}$. Once positions are ordered by score, positions are closed in order of score from highest to lowest, until the deficit is fully covered. We will denote this strategy as $\pi_{Q}$ and define it via the algorithm that implements it.
Given a haircut budget $B=\theta_{\pi_{Q}}D_{T}(\mathcal{P}_{n})$, the algorithm to choose $h_{i}$ (or equivalently, seized amounts $x_{i}=h_{i}w_{i,T}$) works as follows:
@@
    For each position $\mathfrak{p}=(q,c,\ell,t,b)\in\mathcal{P}_{n}$, define the score [^9] We note that technically, many exchanges define the leverage used in the score differently (i.e., instead of the score being linear in $\ell$, some exchanges multiply by $\frac{\ell pq}{pq+c})$. This does not change our main results and mainly adds technical complications to the example. $s_{T}(\mathfrak{p},p^{ref})$. For the PNL-leverage ranking, we have $s_{T}(\mathfrak{p},p^{ref})=\ell\,\frac{\hat{p}_{T}}{p^{ref}}$.